OpenStudy (anonymous):

Verify the Identity. cos(α - β) - cos(α + β) = 2 sin α sin β

OpenStudy (jdoe0001):

well.. just expand both cos(α - β) and cos(α + β) and see what you get

OpenStudy (jdoe0001):

recall \(\bf cos({\color{brown}{ \alpha}} + {\color{blue}{ \beta}})= cos({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}})- sin({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}}) \\ \quad \\ cos({\color{brown}{ \alpha}} - {\color{blue}{ \beta}})= cos({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}}) + sin({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}})\) expand, simplify, be happy, eat ice-cream

OpenStudy (bee_see):